OpenStudy (anonymous):

(a+bi)^2=i What is square(i) solve geometrically in a complex number plane

OpenStudy (anonymous):

squareroot of i

OpenStudy (anonymous):

http://www.milefoot.com/math/complex/squarerootofi.htm this website helps u a lot i think

OpenStudy (helder_edwin):

u can expand the square \[ \large i=(a+ib)^2=a^2+2abi-b^2=(a^2-b^2)+2abi \]

OpenStudy (helder_edwin):

from this \[ \large a^2-b^2=0 \] and \[ \large 2ab=1 \]

OpenStudy (anonymous):

yeah but we need a geometrical solution

OpenStudy (anonymous):

I need to setch how I did it

OpenStudy (anonymous):

or can I just answer squareroot of 2 / ?

OpenStudy (helder_edwin):

well the equations give us \(a=b\) and \(a=1/\sqrt{2}\)

OpenStudy (anonymous):

I dont understand....

OpenStudy (helder_edwin):

what?

OpenStudy (anonymous):

nothing now I see

OpenStudy (anonymous):

Thanks!

OpenStudy (helder_edwin):

but ... i have no idea how to do this "geometrically".

OpenStudy (anonymous):

neither have I T_T

OpenStudy (anonymous):

I can see how to do it algebraic and thats how I will turn it in I guess

OpenStudy (helder_edwin):

maybe if u write the complex number \(z=a+ib\) in trigonometric form. that way u could use a geometric approach.